Funzioni Trigonometriche

Formule importanti:

sen^{2}(α) + cos^{2}(α) = 1, tg(α) = \frac{sen(α)}{cos(α)}.

sen(α + β) = sen(α) cos(α) + cos(α)sen(α); sen(α − β) = sen(α) cos(α) − cos(α)sen(α); cos(α + β) = cos(α) cos(β) − sen(α)sen(β); cos(α − β) = cos(α) cos(β) + sen(α)sen(β).

\begin{cases}sen(2α) = 2sen(α) cos(α); \\ cos(2α) = cos^{2}(α) − sen^{2}(α) = 2 cos^{2}(α) − 1 = 1 − 2sen^{2}(α). \end{cases}

Semplificare le seguenti espressioni: 1-4. cos(π − α), tg(π + α), sen(2π − α), cos(−α); 5-8. tg(π − α), cos(3π − α), sen(π + α), sen(π − α); 9-12. sen(\frac{π}{2} − α), cos(\frac{π}{2} + α), tg(\frac{3π}{2} − α), sen(\frac{3π}{2} + α); 13-16. sen(\frac{π}{2} + α), cos(\frac{π}{2} − α), tg(\frac{3π}{2} + α), sen(\frac{3π}{2} − α); 17-20. 1 − sen^{2}(α), 1 + tg^{2}(α), \frac{1}{tg(α)}, cos^{2}(α) − 1.

Risolvere le seguenti equazioni:

21. sen(x) = −\frac{1}{2} \sqrt{2}, cos(x) = \frac{1}{2} \sqrt{3}, tg(x) = −1, sen(x) = \frac{1}{2};

22. cos(x) = −\frac{1}{2}, cos(x) = \frac{1}{2} \sqrt{2}, tg(x) = −\sqrt{3}, cos(x) = −\frac{1}{2} \sqrt{3};

23. tg(x) = 0, sen(x) = −1, cos(x) = 0, cos(x) = −1;

24. tg(x) = \frac{1}{3} \sqrt{3}, cotg(x) = 1, sen(x) = −\frac{1}{2}, cos(x) = −\frac{1}{2} \sqrt{3}.

Tracciare i grafici delle seguenti funzioni: 25-28. sen(x), cos(x), tg(x), cotg(x); 29-32. sen(2x), cos(3x), tg(x/2), tg(π − x); 33-36. sen(x + \frac{π}{4}), cos(\frac{π}{3} − x), cos(x + π), tg(\frac{π}{4} − x); 37-40. tg(−2x), sen(π − 2x), cos(π + 3x), cos(2x − π).